OpenStudy (anonymous):

express sin 150 degrees as the cosine of a positive acute angle

Parth (parthkohli):

\[\sin(x) = \cos(90 - x)\]\[\sin(150) = \cos(90 - 150) = \cos(-60) = \cos(60)\]

OpenStudy (anonymous):

thank you so much!

Parth (parthkohli):

Hmm, no problem.

OpenStudy (gudden):

ahann! you can solve that question @ParthKohli